OpenStudy (anonymous):

Simplify x squared minus 16 times x plus 64 all over x squared minus 64.. A:−1 B:x minus 1 all over x plus 1 C:x minus 8 all over x plus 8 D:x minus 8 all over x squared minus 64

OpenStudy (anonymous):

\[\frac{x^2 - 16x + 64}{x^2 - 64} = \frac{(x-8)(x-8)}{(x+8)(x-8)} = \frac{x-8}{x+8}\]